OpenStudy (anonymous):

write equations for the lines repersented by pairs of points using point-slope formula : (3,2) & (8,2)

OpenStudy (anonymous):

slope of line=(2-2)/(8-3)=0 equation is (y-2) =0(x-3) , y=2 is the ans